OpenStudy (anonymous):

subtract 5 4/5 - 1 2/3

OpenStudy (anonymous):

62/15

OpenStudy (anonymous):

what s the whole #

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4 62/15

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4 2/15

OpenStudy (anonymous):

okay so for this question subtract 5 4/5 - 1 2/3 the answer is 4 2/15?

OpenStudy (anonymous):

\[\left(5+\frac{4}{5}\right)-\left(1+\frac{2}{3}\right)=\frac{62}{15}=4\frac{2}{15} \]

OpenStudy (anonymous):

I have another question for you please Add 5 2/25 + 2 2/15

OpenStudy (anonymous):

\[7\frac{16}{75} \]